OpenStudy (anonymous):

Factor the polynomial 81x^4−16y^4

OpenStudy (anonymous):

@primeralph

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

Use the identity: a^2 - b^2 = (a + b)(a - b)

OpenStudy (anonymous):

(9-2)(9+2)

OpenStudy (ranga):

81x^4−16y^4 = (9x^2)^2 - (4y^2)^2 = (9x^2 + 4y^2)(9x^2 - 4y^2) You can further factor (9x^2 - 4y^2) using the exact same method again.

OpenStudy (anonymous):

(3x-2y) (3x+2y)??

OpenStudy (ranga):

Yes.

OpenStudy (anonymous):

2x^3+16y^3?

OpenStudy (anonymous):

new question ^^

OpenStudy (ranga):

Factor out 2 first. Then use the identity: a^3 + b^3 = (a + b)(a^2 - ab + b^2)

OpenStudy (anonymous):

2(x^2+8y^2)

OpenStudy (ranga):

2x^3+16y^3 = 2(x^3 + 8y^3) = 2{ (x)^3 + (2y)^3 } = ? Use the above identity for the terms inside { }

OpenStudy (anonymous):

wont the exponent be 2 instead of 3?

OpenStudy (ranga):

No. You are just factoring out a number 2 first.

OpenStudy (anonymous):

can you show me step by step for to get the answer?

OpenStudy (ranga):

2x^3+16y^3 = 2(x^3 + 8y^3) = 2{ (x)^3 + (2y)^3 }. Use the identity: a^3 + b^3 = (a + b)(a^2 - ab + b^2) 2{ (x)^3 + (2y)^3 } = 2{ (x + 2y)( x^2 - (x)(2y) + (2y)^2 ) } = 2(x + 2y)(x^2 - 2xy + 4y^2)

OpenStudy (anonymous):

x^3z^6−27y^9 @ranga

OpenStudy (anonymous):

step by step please

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

It is using the same principle again. Make use of the identity: a^3 - b^3 = (a - b)(a^2 + ab + b^2)

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

can you go step by step again?

OpenStudy (ranga):

I will do the first step. You should be able to do the remaining steps. x^3z^6−27y^9 = (xz^2)^3 - (3y^3)^3 Make use of the identity: a^3 - b^3 = (a - b)(a^2 + ab + b^2) Here a = xz^2 and b = 3y^3. Just put a and b into the right hand side.

OpenStudy (anonymous):

Im confused

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

x^3z^6−27y^9 = (xz^2)^3 - (3y^3)^3 = (xz^2 - 3y^3)(x^2z^4 - 3xy^3z^2 + 9y^6)